### v1

# Kang Zhu1,2, Qianbo Zang911footnotemark: 1, Shian Jia411footnotemark: 1, Siwei Wu1,311footnotemark: 1, Feiteng Fang6,1011footnotemark: 1 Yizhi Li1,3,Shuyue Guo1, Tianyu Zheng1,2, Bo Li5, Haoning Wu2, Xingwei Qu1

Abstract
With the remarkable success achieved by Multimodal Large Language Models (MLLMs), numerous benchmarks have been designed to assess MLLMs’ ability to guide their development in image perception tasks (e.g., image captioning and visual question answering). However, the existence of numerous benchmarks results in a substantial computational burden when evaluating model performance across all of them. Moreover, these benchmarks contain many overly simple problems or challenging samples, which do not effectively differentiate the capabilities among various MLLMs. To address these challenges, we propose a pipeline to process the existing benchmarks, which consists of two modules (i.e., Semi-Automated Screening Process and Eliminating Answer Leakage). The Semi-Automated Screening Process filters out samples that cannot distinguish the model’s capabilities by synthesizing various MLLMs and manually evaluating them. The Eliminate Answer Leakage filters the samples whose answers can be inferred without images. Finally, we curate the LIME-M: Less Is More for Evaluation of Multimodal LLMs, which is a lightweight Multimodal benchmark and can more effectively evaluate the performance of different models. Our experiments demonstrate that 1. LIME-M can better distinguish the performance of different MLLMs within less sample numbers (24% of original) and time spent (23% of original); 2. LIME-M eliminates answer leakage and mainly focuses on the information within images; 3. The current automatic metric (i.e., CIDEr) is insufficient for evaluating MLLMs’ capabilities in captioning. Moreover, we find that removing the caption task score when calculating the overall score demonstrates a more precise reflection of model performance differences. All our codes and data are released at https://github.com/kangreen0210/LIME-M
1 Introduction
Figure 1: Pipeline of the Data Curation. The left half part is the Open-Source Models as Judges module, which uses several Multimodal LLMs to answer questions for each sample and assess their difficulty. The upper right part is the Semi-Automated Screening Process module filtering some samples that are too simple or difficult. As for the Eliminating Answer Leakage, we filter the sample that can be answered without the image.
In order to better understand the model’s capabilities and guide to address the shortcomings of MLLMs, researchers have developed numerous benchmarks for various tasks (Antol et al., 2015; Wei et al., 2023; Fu et al., 2023; Yue et al., 2024; Wu et al., 2024a). These benchmarks thoroughly explore the capabilities of MLLMs in various tasks such as image captioning, image question answering, and Multimodal retrieving.
However, existing MLLM benchmarks and unified evaluation frameworks cannot effectively and efficiently reflect the ability of MLLMs. Current benchmarks include numerous relatively simple samples (i.e., how many chairs are in the picture) and some incorrect questions caused by annotation issues. Most MLLMs will consistently perform on these samples (i.e., all correct or all wrong). Therefore, those benchmarks cannot fully reflect the gap between different MLLMs and across various tasks. Besides, the current unified Multimodal evaluation frameworks require significant computational resources, necessitating integrating much evaluation data from various benchmarks. How to select effective evaluation data has been largely overlooked by current researchers.
To address the aforementioned issues, we propose to use a general data process pipeline and curate a LIME-M, which contains 9403 samples and is refined across 10 tasks within 6 domains. We select six major tasks in the Multimodal domain and use 9 MLLMs to refine those 10 benchmarks within the corresponding domain. To eliminate bias introduced by individual models, we chose 9 models as judges and filter samples based on their performance. On the one hand, we removed samples that most models answered correctly due to the fact that they cannot distinguish the capabilities among different models. On the other hand, we use a method that combines humans and MLLMs to filter out some abnormally difficult samples. Meanwhile, we use LLMs to filter out samples that could be answered directly from the context of the question. After that, we obtain a smaller yet higher-quality unified bench (i.e., LIME-M).
We conducted various experiments on LIME-M using both MLLMs and LLMs on different input settings, such as QA + image inputs, QA input (text-only input), and the QA + image description experiment. We have made several valuable findings:
  • LIME-M can better reflect the performance differences of MLLMs. On our LIME-M benchmark, under consistent conditions (same model series, same model size), different MLLMs demonstrate a wider score range, indicating that LIME-M is more effective at reflecting performance differences between models with a smaller amount of data.
  • MLLMs exhibit varying capabilities across different subtasks. Specifically, they excel in the Visual Question Answering (VQA) subtasks, showcasing relatively high performance when answering questions directly related to factual information depicted in images. However, their performance is comparatively lower in tasks that necessitate the application of additional commonsense knowledge or complex reasoning. This highlights the significant image content recognition capabilities of current MLLMs. Additionally, most MLLMs perform exceptionally well in POPE and Caption subtasks, while demonstrating relatively poor performance in OCR subtasks. This underscores the requirement for MLLMs to perceive deeper content within an image for OCR tasks.
  • Through the correlation analysis of scores across different tasks, we find that using traditional automatic metrics for the captioning task makes it difficult to reasonably evaluate the model’s performance. Different tasks have varying requirements for factual perception and the application of additional commonsense knowledge in images.
2 Method
Most benchmarks contain low-quality, noisy data.
Figure 2 shows the statistics of different subtasks within our LIME-M benchmark. It is worth mentioning that the proportion of easy and wrong samples exceeds 30%. Out of the 10 subtasks, 6 have proportions exceeding 50%. Notably, for the POPE dataset, 95% of the data can be classified as noisy or erroneous. This indicates that existing benchmarks are filled with a large amount of low-quality data, which does not accurately reflect the true capabilities of MLLMs.
To address those issues, we propose a general data curation pipeline, which consists of three main stages: (1) Open-source models as judges, (2) Semi-automated screening process, and (3) Eliminating Answer Leakage. Our approach aims to enhance existing benchmarks by removing inaccurate and oversimplified data.
Figure 2: Overall data statics about selected subtasks. Easy: questions that most models can answer correctly, Bad Case: questions that may contain errors, Remained: questions that finally remain.
2.1 open-source models as judges
To avoid potential biases that may exist in individual MLLMs, we select ten different types of open-source models as judges. To categorize the difficulty of each sample, we analyze the performance of all judge models on each question and label the difficulty based on the number of models that answered correctly. If $N\geq 6$, the question is classified as the easy set. If $3\leq N\leq 5$, it is classified as the middle set. Conversely, if $N\leq 2$, it is classified as the hard set.
2.2 Semi-automated screening process
Easy samples do not effectively differentiate the capabilities of various models, as most models can answer them correctly. Therefore, we remove the easy samples to assess model performance better.
Furthermore, we find that some questions are not correctly answered by any model, which can be due to potential errors in the question design. To mitigate these potential errors and filter out totally incorrect questions, we implement a semi-automated screening process, which consists of two stages. In the first stage, all questions with zero passes are reviewed by GPT-4V to assess their correctness in terms of logic and meaning. In the second stage, questions deemed correct by GPT-4V are then manually screened. This strategy helps us eliminate meaningless or erroneous data from the dataset, thereby reducing its size and improving its quality.
2.3 Eliminating Answer Leakage
Although the previous two stages have filtered out potential errors and assessed the quality of the questions, we still need to address the potential issue of ANSWER LEAKAGE. Multimodal Answer Leakage can be summarized into two main categories: 1.Text Answerable Questions: The textual information contains all the necessary details to answer the question, making the corresponding visual information redundant. 2.Seen Questions: The model has encountered a specific question during training and has memorized the question along with its corresponding ground truth. As a result, some questions can be answered by pure text-based language models (LLMs) without relying on visual information. Therefore, we conduct a text-only check using pure text LLMs. Based on LLMs’ responses, we will remove the samples that can be directly answered without using the image. After that, we proportionally sample 1200 samples from these categories based on their difficulty levels, For benchmarks with fewer than 1,200 entries, we adapt all samples.
3 LIME-M: A comprehensive MLLMs benchmark
Task Domain Dataset Split Full Size Lite Size
Captioning TextCaps COCO-Caption val val 3166 5000 1200 1200
T/F reasoning POPE val 9000 443
Normal VQA OK-VQA TextVQA val val 5046 5000 1200 1200
Infographic QA infoVQA ChartQA val val 2801 2500 1200 1200
Science QA ScienceQA AI2D val val 2097 3088 300 1000
OCR OCRBench val 1000 460
Table 1: Data statics: Full Size: the size of the original dataset, Lite Size: the final size of the LIME-V. For the COCO-Caption dataset, we selected the 2017 subset, and for the ScienceQA dataset, we chose the ScienceQA-IMG subset.
In this section, we propose LIME-M, a comprehensive benchmark for Multimodal Large Language Models (MLLMs). LIME-M streamlines existing mainstream benchmarks. Table 1 shows the main datasets included in our benchmark, as well as the data scale after careful pruning. For each sub-dataset, we aim to keep the size around 1k samples.
3.1 Task definition
We have categorized the existing mainstream tasks into six domains: Captioning, T/F Reasoning, Normal VQA, Infographic Understanding QA, Science QA, and OCR. Below are the task definitions for each domain
  • Image understanding and captioning: The Captioning task focuses on the fundamental image-text understanding ability, requiring MLLMs to accurately describe and understand the content of images. This ability is commonly learned by most multimodal models during the pre-training stage. For example, the CLIP model aligns image and text features through contrastive learning, making Captioning a measure of the basic capabilities of MLLMs.
  • T/F reasoning: T/F Reasoning requires the model to judge the truthfulness of textual statements based on the image content. This not only demands basic image understanding from the MLLMs but also requires a certain level of reasoning ability.
  • Normal VQA: Normal VQA, or Visual Question Answering, comprehensively evaluates the model’s ability to answer questions based on visual input. MLLMs are required to select the most appropriate answer from specific options.
  • Infographic Understanding QA: This task differs from Normal VQA as it tests the model’s ability to retrieve details from images. MLLMs need to find the most relevant information in the image related to the question and then provide a reasoned answer.
  • Science QA: Science QA includes questions and answers related to natural science knowledge. This requires the model to have domain-specific knowledge in natural sciences, mainly assessing the MLLMs’ mastery of knowledge within a specific domain.
  • OCR: The OCR task requires the precise extraction of textual content from images.
3.2 data statistics
LIME-M is composed of 10 open-source multimodal evaluation benchmarks, with scales ranging from 1,000 to 9,000. After our three-stage data curation, the data scale of each benchmark is significantly reduced. Figure 1 shows the number of samples removed at each stage compared to the original dataset. The amount of data removed varies at each stage, with the most being removed in the first stage, reflecting a large number of low-difficulty or data-leakage samples in the existing 9 MLLMs. Comparing the data volumes before and after the second stage of semi-automated screening, we can see that many datasets, such as OK-VQA and TextVQA, have a high rate of low-quality data leading to MLLMs’ incorrect answers. Additionally, some datasets, such as ScienceQA and AI2D, have a significant amount of data removed after the third stage, indicating that many questions in these datasets can be answered correctly by MLLMs without relying on images. The statistics of the curated data are shown in  1. LIME-M includes all data after the third stage of screening, and to further streamline the dataset, we have selected a subset from each screened benchmark based on difficulty, with each subset maintaining a data volume of around 1k.
Figure 3: The number of samples removed at each stage compared to the original data, including three stages of filtering and the final sampling stage.
4 Experiment
4.1 Experiment Setting
To evaluate the quality of LIME-M, we conducted a series of experiments across various open-source and closed-source models. These experiments primarily encompass the following four settings:
  • Main experiment: To demonstrate the performance of LIME-M, mainstream open-source models are evaluated using a standardized process to reflect their overall performance differences.
  • Text only set: To prevent potential data leakage issues, we conducted validation experiments using text-only QA pairs. This was to verify whether LLMs could correctly answer questions based on text-only information.
  • Text-only question with Image Description set: Image Description (ID) refers to simple descriptions of images that represent superficial information contained within them. For most MLLMs, questions containing only superficial information are easy to answer; inversely, questions requiring complex visual inference are significantly more challenging. To further validate whether LIME-M can reflect the capabilities of MLLMs, we input text-only QA pairs combined with ID into LLMs and test their ability.
4.2 Baselines
We select LLaVA-1.5 (Liu et al., 2023a; b), LLaVA-1.6 (Liu et al., 2024), Tinny-LLaVA (Zhou et al., 2024), MiniCPM (Hu et al., 2024), Idefics-2  [^1] https://huggingface.co/blog/idefics2, Deepseek-VL(Lu et al., 2024), CogVLM (Wang et al., 2023; Hong et al., 2023), XComposer-4KHD (Zhang et al., 2023), Mantis (Jiang et al., 2024), InternVL-1.5 and InternVL-2 (Chen et al., 2023; 2024) as our MLLMs baseline, and LLaMA3, Yi, Yi-1.5 (AI et al., 2024), Qwen-1.5 (Bai et al., 2023a) and Qwen2 (Yang et al., 2024) as LLMs baseline. To ensure fairness in the evaluations, we use the unified evaluation framework provided by MLLMs-eval to conduct evaluation experiments on LIME-M. For models not supported by this framework, we refine the inference code provided by the model developers to make it compatible with the new models for the sake of aligning the results of different models.
Metrics
For most tasks included in LIME-M, we reference the metrics computation methods used in MLLMs-eval. Specifically, for tasks such as AI2D, ScienceQA, OCRBench, and POPE, we calculate the accuracy of the extracted responses. For tasks such as OK-VQA and TextVQA, the metric scores are calculated based on the overlap between the response and the candidate answers. For tasks like TextCaps and COCO-Caption2017, we use the CIDEr as the score. The ANLS metric is used for the infoVQA task, and the Relaxed Overall metric is used for the ChartQA task.
We calculate the sub-scores for each task category by taking a weighted average of the subtask scores, and then compute the overall score by weighted averaging the scores of all tasks except for the caption tasks. The details of the metrics calculation can be found in the appendix.
5 Results
5.1 Main Result
T/F Common VQA InfoVQA ScienceQA OCR Captioning
Model Size Overall POPE $\uparrow$ TVQA $\uparrow$ OK VQA $\uparrow$ ChQA $\uparrow$ IVQA $\uparrow$ AI2D $\uparrow$ SciQA $\uparrow$ OCRBen $\uparrow$ COCO $\uparrow$ TCaps $\uparrow$
InternVL-2 2023 ( - ) 40B 64.19 51.69 77.98 19.45 88.33 73.92 69.20 98.33 34.90 63.10 48.94
Qwen2-VL 2023b ($\uparrow$ 1) 7B 62.56 53.05 74.56 28.18 83.17 78.50 58.20 75.00 35.70 68.74 110.60
InternVL-1.5 2024 ($\downarrow$ 1) 26B 62.00 51.69 69.88 35.47 87.00 70.92 54.81 91.67 27.90 69.24 112.63
InternVL-2 2023 ( - ) 26B 61.48 54.63 75.20 16.08 87.67 70.58 62.80 96.33 32.70 76.18 56.91
LLaVA-OneVision 2024 ( - ) 7B 60.28 52.37 65.22 37.32 80.83 68.17 59.20 92.00 22.00 104.74 108.18
InternVL-2 2023 ( - ) 8B 59.52 49.21 66.10 24.20 83.75 68.25 60.40 95.67 32.50 42.55 25.44
XComposer2-4KHD 2023 ($\uparrow$ 4) 7B 55.67 46.28 60.30 28.13 80.42 66.17 54.90 70.00 24.40 97.07 78.07
InternVL-2 2023($\downarrow$ 1) 4B 54.86 47.18 62.29 17.48 81.92 60.50 54.00 94.33 30.90 35.99 21.67
CogVLM-2 2024 ($\uparrow$ 6) 19B 52.02 51.02 69.46 4.92 80.33 59.50 45.00 84.00 31.40 23.67 34.01
Qwen2-VL 2023b ($\uparrow$ 5) 2B 51.59 50.79 70.70 16.87 67.50 65.00 42.90 58.00 31.30 75.06 101.72
InternVL-2 2023 ($\downarrow$ 2) 2B 51.25 50.79 59.56 21.87 71.75 54.00 45.80 92.33 31.00 51.95 42.59
CogVLM-1 2023 ($\uparrow$ 1) 17B 49.53 55.10 71.20 31.70 61.67 57.25 31.40 53.67 19.30 29.28 38.56
Cambrian 2024 ($\downarrow$ 5) 34B 48.78 49.44 57.28 22.03 71.83 43.17 54.80 78.33 18.00 4.27 4.97
Cambrian 2024 ($\downarrow$ 4) 13B 47.06 50.79 58.93 24.13 69.25 42.83 45.20 62.67 19.50 7.40 6.52
Cambrian 2024 ($\downarrow$ 3) 8B 46.41 49.89 59.00 25.23 69.42 37.67 43.60 69.00 19.80 5.85 6.41
InternVL-2 2023 ($\uparrow$ 2) 1B 45.88 52.82 57.55 15.37 65.83 46.25 37.00 84.33 29.90 15.19 13.19
LLaVA-1.6 2024 ($\uparrow$ 3) 34B 42.73 47.00 51.20 10.40 64.33 42.08 49.60 64.67 17.10 84.25 48.25
MiniCPM-LLaMA3-2.5 2024 ($\downarrow$ 3) 8B 40.62 43.10 61.80 25.30 69.00 48.10 6.60 6.60 25.70 31.91 39.86
LLaVA-OneVision 2024 ($\uparrow$ 4) 0.5B 39.85 48.98 48.61 23.13 55.42 40.67 31.20 53.00 19.70 96.40 90.28
LLaVA-LLaMA3 2023 ($\downarrow$ 3) 8B 39.81 44.24 40.01 34.72 64.42 22.25 40.72 61.67 13.90 99.35 48.71
Mantis-Idefics-2 2024 ( - ) 8B 38.09 44.24 44.51 29.07 59.00 20.50 35.80 70.00 14.80 61.82 103.07
Deepseek-VL 2024 ($\uparrow$ 2) 7B 37.17 48.50 44.80 25.00 54.67 22.33 37.00 68.33 11.70 54.22 83.21
LLaVA-1.6-vicuna 2024 ($\downarrow$ 4) 13B 35.93 43.10 43.90 16.10 54.50 28.75 38.10 53.00 14.70 76.62 47.83
Idefics-2 2024 ($\downarrow$ 2) 8B 34.85 42.00 56.50 35.60 13.08 23.92 38.10 78.67 19.60 61.23 94.51
LLaVA-1.6-vicuna 2024 ( - ) 7B 29.27 41.10 39.00 12.50 43.08 22.67 27.10 47.33 10.90 76.05 48.35
Mantis-SigLIP 2024 ($\uparrow$ 1) 8B 28.70 45.60 26.34 32.43 35.33 16.25 27.70 62.67 4.90 68.16 81.21
MiniCPM 2024 ($\uparrow$ 2) 1.0 25.58 44.00 37.00 6.20 35.75 13.42 27.90 60.67 6.70 68.96 76.65
LLaVA-1.5 2023a ($\downarrow$ 2) 13B 20.12 36.60 19.50 32.10 5.50 12.42 25.90 48.33 2.70 80.89 68.73
LLaVA-1.5 2023a ($\downarrow$ 1) 7B 16.98 32.51 16.50 23.43 5.25 9.08 24.05 49.00 2.20 79.20 65.73
InstructBLIP-vicuna 2023 ( - ) 7B 15.30 45.10 11.40 22.10 3.00 9.08 21.90 34.33 2.00 102.08 53.14
Tiny-LLaVA-1 2024 ( - ) 1.4B 13.76 37.00 18.70 0.90 4.50 12.17 22.80 44.67 1.60 63.19 58.91
InternVL-2 40B 80.31 89.23 82.59 50.98 85.52 76.08 85.88 98.56 79.90 99.15 62.03
InternVL-1.5 26B 79.49 88.90 79.00 60.70 83.70 72.50 78.90 94.50 71.40 95.80 148.10
Qwen2-VL 7B 79.14 88.17 80.92 55.68 83.32 78.86 80.73 85.57 81.20 92.13 144.36
InternVL-2 26B 78.82 88.64 82.06 48.50 84.44 72.72 83.16 97.47 77.60 110.30 80.10
LLaVA-OneVision 7B 78.71 89.17 76.02 60.98 80.12 70.69 81.38 95.88 62.10 140.45 136.97
InternVL-2 8B 77.84 87.90 77.00 52.02 82.48 70.65 82.25 97.03 76.50 89.77 36.70
InternVL-2 4B 73.97 87.71 74.51 38.43 81.04 65.19 78.08 96.03 75.00 54.08 30.17
Cambrian 34B 73.26 88.46 72.11 52.07 74.60 51.48 80.41 85.52 59.00 8.18 6.08
InternVL-2 2B 73.00 88.90 72.39 43.74 74.72 57.69 72.70 94.25 75.50 79.52 59.81
Cambrian 13B 72.39 88.53 73.07 53.28 72.60 50.73 73.93 79.08 61.40 14.33 9.44
XComposer-4KHD 7B 71.93 87.00 74.30 51.90 80.60 72.80 34.40 96.00 66.90 134.00 111.40
Cambrian 8B 71.84 88.24 72.47 52.17 73.44 48.05 72.99 80.32 61.60 9.13 7.97
CogVLM-1 17B 71.34 88.90 79.70 46.90 67.00 63.30 61.90 70.50 59.10 28.40 44.70
MiniCPM-LLaMA3-2.5 8B 71.22 88.00 75.00 52.30 72.90 56.90 71.70 53.00 69.40 35.50 52.90
Qwen2-VL 2B 70.86 87.78 78.70 40.59 73.12 67.15 70.21 77.89 75.30 103.52 131.92
CogVLM-2 19B 69.93 87.56 77.59 18.51 79.84 62.62 72.41 90.93 76.60 24.10 42.23
LLaVA-LLaMA3 8B 69.74 87.80 65.40 60.20 69.30 37.60 71.60 73.30 55.00 135.00 69.60
InternVL-2 1B 68.46 87.94 69.67 33.84 71.40 52.02 62.56 89.59 74.20 49.34 18.03
LLaVA-1.6-vicuna 13B 67.29 87.50 67.00 46.30 62.20 41.50 70.40 73.50 55.00 101.90 67.30
LLaVA-1.6 34B 67.22 85.60 68.90 31.00 67.40 51.90 76.10 82.70 58.60 114.40 69.10
Mantis-Idefics-2 8B 66.91 86.90 63.51 52.50 63.56 31.17 66.81 81.80 54.20 79.42 134.08
Idefics-2 8B 66.73 86.80 71.30 53.90 26.40 37.00 69.20 87.20 61.60 71.90 119.10
LLaVA-OneVision 0.5B 65.65 88.33 65.85 44.17 61.36 46.23 57.09 67.03 57.60 131.90 120.81
Deepseek-VL 7B 65.62 87.10 63.20 48.70 60.60 34.30 63.40 81.70 43.30 67.60 110.10
LLaVA-1.6-vicuna 7B 64.80 87.60 64.90 44.20 55.00 37.00 65.30 70.20 52.40 100.00 72.00
LLaVA-1.5 13B 59.58 87.10 48.70 58.30 18.10 29.50 59.40 72.80 33.60 115.40 104.00
Mantis-SigLIP 8B 58.96 81.47 49.59 52.90 42.56 26.56 57.84 75.36 34.50 91.37 111.43
LLaVA-1.5 7B 57.27 87.00 46.10 53.40 18.20 25.80 55.20 69.50 31.50 109.00 98.00
MiniCPM 1.0 56.18 85.10 55.30 47.30 15.40 20.10 56.90 43.00 60.00 25.90 41.60
InstructBLIP-vicuna 7B 47.87 85.00 33.20 45.20 12.50 22.90 34.00 36.40 25.90 141.40 74.00
Tiny-LLaVA-1 1.4B 34.30 56.30 38.50 3.80 11.10 22.20 32.30 58.20 17.20 80.90 83.10
Table 2: Comparing overall scores of LIME-M and Original. Top: results on LIME-M, Bottom: results on the original dataset. The arrow next to the model name indicates the change in ranking on LIME-M compared to the original dataset. $\uparrow$: upward shift, $\downarrow$: downward shift, and -: no change.
As shown in Tab 2, we evaluate the mainstream MLLMs on our LIME-M benchmark. Overall, models with larger parameter sizes and newer versions tend to have higher overall scores. The InternVL-1.5, InternVL-2-Large (26B,40B) and LLaVA-OneVision-7B achieve the best overall performance, whose overall scores all surpass 60%. The performance of InternVL-2-Small(1B-8B), CogVLM series, and Cambrian series are second only to InternVL-2-Large and LLaVA-OneVision-7B, with their overall scores ranging from 45% to 60%. Comparing the overall score of LIME-M and Origin benchmarks, it can be observed that certain model series, such as Cambrian and LLaVA-1.5, their overall scores experience a decline. Conversely, the CogVLM and LLaVA-OneVision series’ overall scores show an improvement. Specifically, that of CogVLM2 and XComposer-4KHD experience significant fluctuations, improved by 4% and 6% respectively
As for the OCR task, all those MLLMs performance are all below 35%, and the performance of most models falls between 10% and 30%. This indicates that OCR tasks remain challenging for current MLLMs. Regarding with POPE subtask and caption subtask, most models demonstrate good performance. These tasks involve generating or assessing descriptions of the content in images, which proves that current MLLMs have strong image content recognition capabilities. As for the VQA task, current MLLMs have a relative high performance on the TextVQA, ChatQA and ScienceQA, which design questions directly ask the facts in the picture. While their performances are relative less on OK-VQA, infoVQA and AI2D, which require to use extra commonsense knowledge or complex reasoning to answer question. This demonstrates that the current MLLMs exhibit significant image content recognition capabilities. However, their ability to perform complex reasoning by utilizing additional knowledge is not strong. We believe this limitation may be due to the constraints of the language model component of MLLMs.
Figure 4: Correlation distribution between LIME-M and Wildvison Elo.
5.2 Correlation Analysis
Figure 4 illustrates the correlation between the various sub-tasks in LIME-M and WildVision Bench. Most tasks in LIME-M exhibit a strong positive correlation with WildVision Bench. There are six subtasks that have correlation scores exceeding 80%. Additionally, the overall score of LIME-M correlates at 91% with WV-Elo, which is higher than any individual sub-task and the original bench’s correlates, demonstrating that the overall score of LIME-M provides a more comprehensive reflection of MLLMs capabilities.
Automated evaluation metrics (e.g., CIDEr) cannot effectively assess the performance of LVMs in captioning tasks.
As an earliest foundational problem, the captioning task has been extensively studied and MLLMs have an exceptional ability on this task. For instance, earlier models like InstructBlip performed exceptionally well on captioning tasks, and there is a broad presence of training data for image captioning in MLLMs’ training processes. However, the captioning task shows a negative correlation with all other sub-tasks. This indicates that previous metrics (e.g., BLEU, CIDEr) only focus on the overlap between the model-generated responses and the ground truth, but do not consider that MLLMs might generate content that is semantically close to the ground truth (i.e., the model-generated response may be semantically similar to the ground truth but expressed differently, or the model may generate more detailed content about the image). Consequently, we exclude it from the overall score calculation.
There is a certain degree of correlation between the sub-tasks in LIME-M.
On the one hand, the relevance of TextVQA, InfographicVQA and OCRBench are relatively high. As shown in the Tab 4, the correlation of those tasks all surpass 85%, and these two VQA tasks require MLLMs to understand the fine-grained content in images to answer question. It demonstrates that OCR tasks also rely on the ability of MLLMs to perceive fine-grained objective facts in images. On the other hand, POPE, ChartQA and InfographicVQA all need the reasoning ability using extra commonsense knowledge. The correlation scores of these tasks are all above 70%, and the POPE requires model to use the extra knowledge to solve the hallucination of MLLMs. We assume that ChartQA and infoVQA may also necessitate the use of additional common knowledge by the models to solve problems.
5.3 Effectiveness of LIME-M
Figure 5: with the same series of models, the distribution differences of various Parameter sizes. Left($\bigstar$): LLaVA-1.6 series, Right($\blacktriangle$): InternVL-2 series
Model series Dataset GiNi stdev
InternVL-2 LIME-M Original 0.063 0.030 6.864 4.421
Cambrian LIME-M Original 0.006 0.002 1.227 0.715
LLaVA-1.6 LIME-M Original 0.042 0.004 6.730 1.418
Table 3: Statistics on the score distributions across different model series
Model size Dataset GiNi stdev
7B LIME-M Original 0.271 0.086 19.041 10.836
8B LIME-M Original 0.128 0.046 10.685 6.270
13B LIME-M Original 0.174 0.043 13.536 6.446
Table 4: Statistics on the score distributions across different model sizes.
LIME-M provides a more challenging evaluation for MLLMs.
As shown in Tab 2, The MLLMs’ performances on the LIME-M are less than that on Original Bench on most tasks. Compared to the Origin benchmark, different MLLMs show a larger score range on our LIME-M, indicating that our LIME-M can better reflect the performance differences between models with a smaller amount of data.
Furthermore, we compare the score variations across different model series and model sizes. Figure 5 illustrates a clear positive correlation between model performance and model size within the same model series. Notably, LIME-M exhibits a more dispersed score distribution, effectively highlighting the differences in model performance. In Table 5.3 and 5.3, the Gini coefficient and standard deviation are used to measure the differences in overall score distribution across same model series and model sizes. The larger the Gini coefficient and standard deviation, the greater the disparity in data distribution. It can be observed that, whether within the same model series or the same model size, LIME-M achieves higher Gini and standard deviation values compared to the original bench. This indicates that LIME-M can better differentiate the performance differences between various models.
AI2D ScienceQA
Model LIME-M Original LIME-M Original
LLaMA3-8B 18.10 46.76 33.33 59.35
LLaMA3-70B 25.70 62.05 56.00 69.91
Qwen1.5-32B 24.10 61.14 43.67 67.97
Qwen1.5-72B 19.80 57.45 35.00 61.13
Qwen2-7B 21.00 57.09 43.00 67.38
Qwen2-72B 20.60 69.95 38.67 63.36
Yi-1.5-9B-Chat 20.10 23.22 17.33 23.60
Yi-1.5-34B-Chat 23.60 54.15 42.00 65.20
Yi-1.5-34B-Chat 25.20 60.69 46.00 70.55
Figure 6: Comparing text only results of LIME-M and original bench. Left: text only results between LIME-M and Original on AI2D and ScienceQA; Right: average score comparison of Original and LIME-M.
LIME-M eliminates potential data leakage.
For multimodal question answering tasks, visual information input is essential, and LLMs are unable to provide correct answers due to they cannot perceive the content within the image. However, as shown in Figure 6 (right), there are severe data leakage issues in the original Bench for the AI2D and ScienceQA tasks. The average score for AI2D is close to 55%, and for ScienceQA, it exceeds 60%, which shows that data from AI2D and ScienceQA in Original are highly likely to have been exposed to the training data of LLMs. In contrast, the LIME-M has eliminated this potential threat, achieving scores below 25% in AI2D and close to 40% in ScienceQA.
According to the results in Figure 6 (left), LLMs tend to score higher on ScienceQA than on AI2D. This is because the former includes many commonsense text-only questions, such as ”What is the capital of Massachusetts?” which are unrelated to the images.
5.4 The Impact of Detail Image Perception
Setting Models AI2D ChQA COCO IVQA OCRBen OK VQA POPE SciQA TCaps TVQA
LLaMA3-8B 23.5 6.4 2.8 12.9 4.2 17.4 32.1 16.4 5.3 17.9
LLaMA3-70B 24.0 7.7 3.3 12.3 4.3 21.8 38.1 39.4 6.0 22.0
Qwen1.5-32B 28.8 6.7 6.5 9.4 4.0 4.7 39.3 46.6 9.2 13.7
Qwen1.5-72B 25.4 2.5 3.2 10.1 4.1 7.8 42.7 44.2 6.0 15.2
Qwen2-7B 27.6 6.7 6.9 11.2 4.1 15.0 44.2 45.5 12.5 19.0
Qwen2-72B 26.3 6.9 2.7 10.8 4.4 10.6 36.3 45.2 5.2 16.8
VB Yi-1.5-9B-Chat 22.1 2.3 0.3 3.1 0.0 7.8 40.0 0.0 0.2 5.8
LLaMA3-8B 49.0 11.4 3.1 18.6 19.3 32.5 46.9 59.5 6.5 26.4
LLaMA3-70B 52.0 12.4 3.6 17.6 19.5 36.4 5.2 64.6 7.8 36.2
Qwen1.5-32B 60.5 10.7 8.1 15.0 20.2 15.8 47.4 68.8 10.6 22.1
Qwen1.5-72B 58.8 6.4 3.8 16.6 20.2 21.1 35.1 68.4 7.1 27.4
Qwen2-7B 59.2 12.7 7.4 19.7 19.6 30.5 44.6 69.0 15.4 33.3
Qwen2-72B 60.4 10.5 3.5 15.7 20.5 24.2 34.3 67.9 6.8 28.7
OB Yi-1.5-9B-Chat 24.7 3.1 0.0 5.9 0.5 7.8 32.7 31.7 0.2 5.8
Table 5: Text-only with VD results: With the condition of providing only text QA information and VD information, the performance comparison between vlms-bench and origin bench.
In our data cleaning process, we removed many questions that most models could answer, as well as a small number of questions that are difficult for both humans and GPT-4 to answer, in order to make the benchmark better highlight the differences in model capabilities. As shown in Tab 5, to investigate whether the remaining samples need to be answered by using textual and image information, we conducted experiments using LLMs to generate answers on both the Original Benchmark and MLLMs Benchmark under QVD (question + visual description) setting.
LIME-M requires MLLMs to perceive deeper levels of image information.
Especially in tasks such as AI2D, OCRBench, and TCaps, the scores of LLMs on LIME-M are significantly lower than on the Original Benchmark when provided with only the questions and simple image descriptions. This indicates that, after removing some of the simpler questions, LIME-M is better at testing the models’ ability to perceive image details.
6 Related work
In recent years, there has been increasing attention on establishing evaluation benchmarks to assess the performance of MLLMs in different scenarios to guide the development of MLLMs. Early multimodal evaluation benchmarks primarily focused on single tasks, such as Visual Question Answering (VQA) (Antol et al., 2015; Goyal et al., 2017; Kafle & Kanan, 2017; Singh et al., 2019; Marino et al., 2019), Image Captioning (Agrawal et al., 2019), and Information Retrieval (Wei et al., 2023). As MLLMs develop, simple benchmarks are no longer sufficient to evaluate the versatile capabilities of these models comprehensively, since most MLLMs have exceptional ability on those benchmarks. Consequently, numerous more difficult and diverse evaluation benchmarks have emerged in recent years to assess the capabilities of MLLMs comprehensively. For instance, MMMU (Yue et al., 2024) and CMMMU (Zhang et al., 2024), are comprehensive benchmark tests for university-level multidisciplinary multimodal understanding and reasoning, spanning various disciplines and fields such as Arts and Design, Business, Science, Health Medicine, Humanities and Social Sciences, and Technology Engineering. MMBench (Liu et al., 2023c) has developed a comprehensive evaluation pipeline that offers fine-grained capability assessment and robust evaluation metrics. MMRA (Wu et al., 2024b) systematically establishes an association relation system among images to access the multi-image relation mining ability of MLLMs. However, those benchmarks cannot distinguish the performance gaps among different models excellently, due to they still have some too simple or difficult samples that most models have the same results. Furthermore, training datasets across different models may contain the samples of those benchmarks, which results in data leakage issues (Fu et al., 2023). Therefore, we propose LIME-M: a comprehensive MLLMs benchmark. We have carefully selected six types of tasks from the existing mainstream benchmarks and scaled down the existing benchmarks according to clear guidelines. It is a streamlined version that retains the essential elements of mainstream MLLMs benchmarks.
7 Conclusion
As MLLMs continue to advance, a notable absence of convenient and high-quality multimodal benchmarks has emerged. In response to this, we propose a pipeline aimed at semi-automatically refining existing benchmarks to enhance their quality, culminating in the development of LIME-M, which comprises 9,400 evaluation samples across 6 types of tasks and 10 different benchmark datasets. By refining the original benchmarks to filter question difficulty and eliminate potentially problematic items, LIME-M offers a more rigorous evaluation for MLLMs, necessitating a deeper understanding of image information. The outcomes of our evaluation experiments demonstrate the heightened challenge posed by LIME-M for MLLMs. We anticipate that our approach will contribute to the advancement of MLLM evaluation systems, and we are committed to continually enriching LIME-M with an expanded array of datasets through regular updates and expansions. Our ultimate goal is to provide the community with a simpler, more efficient, and more accurate evaluation method and suite for MLLMs.
Appendix A Appendix
A.1 overall Data statics
Figure 7: The overall percentage distribution of LIME-M.
Figure 8: Overview of data statics.
A.2 Pipeline details
A.2.1 prompt template details
semi-automated screening process prompt
We chose GPT-4V as the basis for automatic judgment and made requests to the GPT-4V API using the prompt format below. For different subtask types, we used different prompt templates.
For VQA tasks, the prompt template is as follows:
Please judge whether the <Answer> is the golden answer to the <Question>.
If it is, please reply YES, otherwise reply NO.
<Question>: {question}
<Answer>: {answer}
<Your judgement>: <YES or NO>
For captioning tasks, the prompt template is as follows:
Now there is an image captioning task.
Please first describe the content of the image, then compare the image content with the provided captions.
If the captions are suitable as captions for the image, please answer YES; if they are not suitable, please answer NO.
Respond with NO if any of the captions are unsuitable. Respond with YES only if all captions are suitable.
You only need to reply with either YES or NO in <Your judgment>.
<Captions>: {answer}
<Description>: <Content of the image>
<Your judgement>: <ONLY YES or NO>
Exact Vision Description Prompt
For the QVD experiment, we use LLaVA-NEXT-110B to extract information from the images, with the following prompt:
[…]
A.2.2 metrics
  • Subtask metrics: As shown in the table 6, different metrics are used for different subtasks. It is important to note that, except for the CIDEr metric, all other metrics have a range between 0 and 1. The final score for each subtask is calculated by taking the average of these metrics.
    Metric Subtask Formula
    Accuracy AI2D, ScienceQA-IMG, OCRBench, POPE $\text{Accuracy}=\begin{cases}1,&\text{if the prediction is correct}\\ 0,&\text{if the prediction is incorrect}\end{cases}$
    CIDEr TextCaps,COCO-Caption $\text{CIDEr}=\frac{1}{m}\sum_{i=1}^{m}\sum_{n=1}^{N}w_{n}\cdot\frac{g_{i}^{(n)}\cdot r_{i}^{(n)}}{\|g_{i}^{(n)}\|\|r_{i}^{(n)}\|}$
    Match score OK-VQA,TextVQA $\text{SCORE}=\min\left(1,\frac{\text{match\_nums}}{3}\right)$
    ANLS InfoVQA $\text{ANLS}(X,Y)=1-\frac{\text{Lev}(X,Y)}{\max(|X|,|Y|)}$
    Relaxed Overall ChartQA $\text{SCORE}=\frac{\left|\text{prediction}-\text{SCORE}\right|}{\left|\text{target}\right|}$
    Table 6: Metrics for different subtask
  • Overall metric: For the overall metric, we explored two mainstream calculation methods: arithmetic mean 1 and weighted mean 2.
    Arithmetic $\text{Arithmetic Mean}=\frac{1}{n}\sum_{i=1}^{n}x_{i}$ (1)
    Weighted $\text{Weighted Mean}=\frac{\sum_{i=1}^{n}w_{i}x_{i}}{\sum_{i=1}^{n}w_{i}}$ (2)
    The arithmetic mean directly calculates the average of each subtask’s scores, while the weighted mean takes into account the number of samples in each subtask. We compare the results of these two calculation methods, as shown in the table 7. Except for Cogvlm, the scores calculated by the weighted average are lower than those by the arithmetic mean, and they show a higher correlation with WV-ELO. This suggests that the weighted average method is slightly superior to the arithmetic mean, as it considers the impact of the number of data points on the overall score, thereby avoiding potential errors caused by uneven data distribution. Therefore, in our work, we ultimately chose the weighted average as the method for calculating the overall score.
    model overall_weighted overall_sum overall_cider WV_bench
    LLaVA-1.6-vicuna-7B 29.27 30.46 36.07 992
    LLaVA-1.6-vicuna-13B 35.93 36.52 41.04 956
    LLaVA-1.6-34B 42.73 43.30 47.12 1059
    CogVLM 49.53 47.66 44.03 1016
    Deepseek-VL 37.17 39.04 43.31 979
    Idefics2 34.85 38.43 43.83 965
    MiniCPM-v-1.0 25.58 28.95 35.79 910
    Tinny-LLaVA-1-hf 13.76 17.79 24.15 879
    LLaVA-1.5-13B 20.12 22.88 32.3 891
    InstructBLIP-vicuna-7B 15.30 18.61 29.56 862
    correlation score 0.91 0.90 0.87 1
    Table 7: Comparison of different overall metrics method
A.2.3 difficulty classification details
For subtasks using the accuracy (acc) metric, where the scores are binary, with only 1 or 0, other tasks may have various possible score distributions (e.g., COCO-Caption, OK-VQA). Therefore, we determine the threshold score based on the overall distribution of subtask scores, and choose the cutoff value that offers the greatest distinction, as shown in table 8, for the metrics ANLS, Relaxed Overall and Accuracy (Acc), the threshold is set to 1.0, for BLEU-4 (for the captioning task, we use the BLEU-4 metric to represent the score for each question), the threshold is set to 0.2, while for Match Score, it is set to 0.6. When the score is greater than the threshold, it is marked as correct; otherwise, it is marked as incorrect.
Metrics bleu4 Match score ANLS Relaxed Overall Acc
Threshold 0.2 0.6 1.0 1.0 1.0
Table 8: Thresholds for Different Metrics
Figure 9: Different query methods. a:single modal query encodes each modality into separate features. b. multi-modal query fuses the feature encodings of two modalities into a merged embedding feature.
A.3 Alignment With Real World User Queries
To bridge the gap between LIME-M and real-world users’ query. We construct a similarity search system between the LIME-M and the real-world benchmarks(such as wildvision bench, and VibeEval) Different from single-modal similarity search systems, our system leverages multimodal information as query instructions, which contains text and image features information, to explore the effectiveness of the methods and construct the optimal similarity search system, we propose the following different approaches:
  • single modal query
    For single-modal query, we use the corresponding modality information for retrieval, such as text-to-text and image-to-image. The CLIP model, which consists of a text encoder and an image encoder (Radford et al., 2021), has achieved excellent performance in representing multimodal information. As shown in Figure 9, We use CLIP’s encoders to encode single-modal features and use cosine similarity as measure metrics. For each data entry in the real-world benchmark, we index the top-k most similar entries sequentially.
  • multi-modal query For multimodal queries, we combine image features and text features as the query. As shown in Figure 9, We adapt both the Clip and Blip models to combine these features. For the clip model, the text and image inputs are processed into separate vector features, which are then combined into a unified vector through a weighted sum. As to the blip model, image features are fused with text features through a cross-attention mechanism. We only keep the top-1 similar data entry based on similarity.
Figure 10: Correlation coefficient between LIME-M-fit and WildVision bench.
In Figure 10, we calculate the correlation coefficient between LIME-M-fit obtained through different query methods and real-world benchmarks. We find that samples retrieved using image queries align better with the WildVision benchmark compared to those retrieved with text queries. In multimodal VQA tasks, text queries often suffer from various forms of repetition, whereas images are unique and provide more significant information.
However, whether using a single-modal or multi-model query method, the Correlation Coefficient obtained by all methods is lower than the baseline score, which suggests that directly querying results from the origin data does not align well with real-world user queries.
Why current multimodal benchmarks can’t reflect real user queries in the wild real world?
Instructions in original benchmark can not address real-world problems. Although there are various multimodal evaluation datasets (TextVQA, infoVQA, ScienceQA, and MMMU) , the instructions in these datasets differs from real-world user instructions. For the same image, the types of instructions can also vary significantly. As illustrated in Figure11, typical VQA tasks require perception and reasoning about detailed content within an image, such as asking ’the population older than 65 in Vietnam in 2020’. In contrast, real-world queries might ask ’how to reproduce the charts using Python’. This discrepancy leads to inherent differences in instructions, even when using retrieval algorithms to find the most relevant dataset questions.
Figure 11: case study about real world query.
Figure 12: A sample bad case of AI2D
Back to List of figures
Figure 13: A sample bad case of AI2D
Back to List of figures
Figure 14: A sample bad case of ChartQA
Back to List of figures
Figure 15: A sample bad case of InfoVQA
Back to List of figures
Figure 16: A sample bad case of InfoVQA
Back to List of figures
Figure 17: A sample bad case of POPE
Back to List of figures
Figure 18: A sample bad case of POPE
Back to List of figures
Figure 19: A sample bad case of POPE
Back to List of figures
Figure 20: A sample bad case of POPE
Back to List of figures
Figure 21: A sample bad case of POPE
Back to List of figures
Figure 22: A sample bad case of OKVQA
Back to List of figures
Figure 23: A sample bad case of OKVQA
Back to List of figures
Figure 24: A sample bad case of OKVQA
Back to List of figures
Figure 25: A sample bad case of TextVQA
Back to List of figures
Figure 26: A sample bad case of TextVQA
Back to List of figures
Figure 27: A sample bad case of TextVQA
Back to List of figures

### v2

@@
-Kang Zhu1,2, Qianbo Zang911footnotemark: 1, Shian Jia411footnotemark: 1, Siwei Wu1,311footnotemark: 1, Feiteng Fang6,1011footnotemark: 1 Yizhi Li1,3,Shuyue Guo1, Tianyu Zheng1,2, Bo Li5, Haoning Wu2, Xingwei Qu1
+King Zhu1,2, Qianbo Zang911footnotemark: 1, Shian Jia411footnotemark: 1, Siwei Wu1,311footnotemark: 1, Feiteng Fang6,811footnotemark: 1, Yizhi Li1,3 Shawn Gavin1, Tuney Zheng1,2, Jiawei Guo1,2, Bo Li5, Haoning Wu2, Xingwei Qu1
 Abstract
-With the remarkable success achieved by Multimodal Large Language Models (MLLMs), numerous benchmarks have been designed to assess MLLMs’ ability to guide their development in image perception tasks (e.g., image captioning and visual question answering).
-However, the existence of numerous benchmarks results in a substantial computational burden when evaluating model performance across all of them.
-Moreover, these benchmarks contain many overly simple problems or challenging samples, which do not effectively differentiate the capabilities among various MLLMs.
-To address these challenges, we propose a pipeline to process the existing benchmarks, which consists of two modules (i.e., Semi-Automated Screening Process and Eliminating Answer Leakage).
-The Semi-Automated Screening Process filters out samples that cannot distinguish the model’s capabilities by synthesizing various MLLMs and manually evaluating them.
-The Eliminate Answer Leakage filters the samples whose answers can be inferred without images.
-Finally, we curate the LIME-M: Less Is More for Evaluation of Multimodal LLMs, which is a lightweight Multimodal benchmark and can more effectively evaluate the performance of different models.
-Our experiments demonstrate that 1.
-LIME-M can better distinguish the performance of different MLLMs within less sample numbers (24% of original) and time spent (23% of original); 2.
-LIME-M eliminates answer leakage and mainly focuses on the information within images; 3.
-The current automatic metric (i.e., CIDEr) is insufficient for evaluating MLLMs’ capabilities in captioning.
-Moreover, we find that removing the caption task score when calculating the overall score demonstrates a more precise reflection of model performance differences.
-All our codes and data are released at https://github.com/kangreen0210/LIME-M
+Multimodal Large Language Models (MLLMs) are measured on numerous benchmarks like image captioning, visual question answering, and reasoning.
+However, these benchmarks often include overly simple or uninformative samples, making it difficult to effectively distinguish the performance of different MLLMs.
+Additionally, evaluating models across many benchmarks creates a significant computational burden.
+To address these issues, we propose LIME (Less Is More for MLLM Evaluation), a refined and efficient benchmark curated using a semi-automated pipeline.
+This pipeline filters out uninformative samples and eliminates answer leakage by focusing on tasks that require image-based understanding.
+Our experiments show that LIME reduces the number of samples by 76% and evaluation time by 77%, while more effectively distinguishing between models.
+Notably, we find that traditional automatic metrics like CIDEr are insufficient for evaluating MLLMs’ captioning performance, and excluding the caption task score yields a more accurate reflection of overall model performance.
+All code and data are available at https://github.com/kangreen0210/LIME.
 1 Introduction
@@
 Besides, the current unified Multimodal evaluation frameworks require significant computational resources, necessitating integrating much evaluation data from various benchmarks.
-How to select effective evaluation data has been largely overlooked by current researchers.
-To address the aforementioned issues, we propose to use a general data process pipeline and curate a LIME-M, which contains 9403 samples and is refined across 10 tasks within 6 domains.
+The selection of effective evaluation data has been largely overlooked by current researchers.
+To address the aforementioned issues, we propose to use a general data process pipeline and curate a LIME, which contains 9403 samples and is refined across 10 tasks within 6 domains.
 We select six major tasks in the Multimodal domain and use 9 MLLMs to refine those 10 benchmarks within the corresponding domain.
@@
 Meanwhile, we use LLMs to filter out samples that could be answered directly from the context of the question.
-After that, we obtain a smaller yet higher-quality unified bench (i.e., LIME-M).
-We conducted various experiments on LIME-M using both MLLMs and LLMs on different input settings, such as QA + image inputs, QA input (text-only input), and the QA + image description experiment.
+After that, we obtain a smaller yet higher-quality unified bench (i.e., LIME).
+We conducted various experiments on LIME using both MLLMs and LLMs on different input settings, such as QA + image inputs, QA input (text-only input), and the QA + image description experiment.
 We have made several valuable findings:
-  • LIME-M can better reflect the performance differences of MLLMs.
-On our LIME-M benchmark, under consistent conditions (same model series, same model size), different MLLMs demonstrate a wider score range, indicating that LIME-M is more effective at reflecting performance differences between models with a smaller amount of data.
+  • LIME can better reflect the performance differences of MLLMs.
+On our LIME benchmark, under consistent conditions (same model series, same model size), different MLLMs demonstrate a wider score range, indicating that LIME is more effective at reflecting performance differences between models with a smaller amount of data.
   • MLLMs exhibit varying capabilities across different subtasks.
@@
 Most benchmarks contain low-quality, noisy data.
-Figure 2 shows the statistics of different subtasks within our LIME-M benchmark.
+Figure 2 shows the statistics of different subtasks within our LIME benchmark.
 It is worth mentioning that the proportion of easy and wrong samples exceeds 30%.
@@
 This indicates that existing benchmarks are filled with a large amount of low-quality data, which does not accurately reflect the true capabilities of MLLMs.
-To address those issues, we propose a general data curation pipeline, which consists of three main stages: (1) Open-source models as judges, (2) Semi-automated screening process, and (3) Eliminating Answer Leakage.
-Our approach aims to enhance existing benchmarks by removing inaccurate and oversimplified data.
+Inspired by MMStar (Chen et al., 2024a), we utilize open-source MLLMs and LLMs as the judges for filtering, specifically, we removing the existing annotation errors.
+The overall pipeline consists of three main stages: (1) Using open-source models as judges, (2) A semi-automated screening process, and (3) Eliminating answer leakage.
+Our approach aims to improve existing benchmarks by removing inaccurate and oversimplified data.
 Figure 2: Overall data statics about selected subtasks. Easy: questions that most models can answer correctly, Bad Case: questions that may contain errors, Remained: questions that finally remain.
@@
 After that, we proportionally sample 1200 samples from these categories based on their difficulty levels, For benchmarks with fewer than 1,200 entries, we adapt all samples.
-3 LIME-M: A comprehensive MLLMs benchmark
+3 LIME: A comprehensive MLLMs benchmark
 Task Domain Dataset Split Full Size Lite Size
@@
 OCR OCRBench val 1000 460
-Table 1: Data statics: Full Size: the size of the original dataset, Lite Size: the final size of the LIME-V.
+Table 1: Data statics: Full Size: the size of the original dataset, Lite Size: the final size of the LIME.
 For the COCO-Caption dataset, we selected the 2017 subset, and for the ScienceQA dataset, we chose the ScienceQA-IMG subset.
-In this section, we propose LIME-M, a comprehensive benchmark for Multimodal Large Language Models (MLLMs).
-LIME-M streamlines existing mainstream benchmarks.
+In this section, we propose LIME, a comprehensive benchmark for Multimodal Large Language Models (MLLMs).
+LIME streamlines existing mainstream benchmarks.
 Table 1 shows the main datasets included in our benchmark, as well as the data scale after careful pruning.
@@
 3.2 data statistics
-LIME-M is composed of 10 open-source multimodal evaluation benchmarks, with scales ranging from 1,000 to 9,000.
+LIME is composed of 10 open-source multimodal evaluation benchmarks, with scales ranging from 1,000 to 9,000.
 After our three-stage data curation, the data scale of each benchmark is significantly reduced.
@@
 The statistics of the curated data are shown in  1.
-LIME-M includes all data after the third stage of screening, and to further streamline the dataset, we have selected a subset from each screened benchmark based on difficulty, with each subset maintaining a data volume of around 1k.
+LIME includes all data after the third stage of screening, and to further streamline the dataset, we have selected a subset from each screened benchmark based on difficulty, with each subset maintaining a data volume of around 1k.
 Figure 3: The number of samples removed at each stage compared to the original data, including three stages of filtering and the final sampling stage.
@@
 4.1 Experiment Setting
-To evaluate the quality of LIME-M, we conducted a series of experiments across various open-source and closed-source models. These experiments primarily encompass the following four settings:
-  • Main experiment: To demonstrate the performance of LIME-M, mainstream open-source models are evaluated using a standardized process to reflect their overall performance differences.
+To evaluate the quality of LIME, we conducted a series of experiments across various open-source and closed-source models. These experiments primarily encompass the following four settings:
+  • Main experiment: To demonstrate the performance of LIME, mainstream open-source models are evaluated using a standardized process to reflect their overall performance differences.
   • Text only set: To prevent potential data leakage issues, we conducted validation experiments using text-only QA pairs.
@@
 For most MLLMs, questions containing only superficial information are easy to answer; inversely, questions requiring complex visual inference are significantly more challenging.
-To further validate whether LIME-M can reflect the capabilities of MLLMs, we input text-only QA pairs combined with ID into LLMs and test their ability.
+To further validate whether LIME can reflect the capabilities of MLLMs, we input text-only QA pairs combined with ID into LLMs and test their ability.
 4.2 Baselines
-We select LLaVA-1.5 (Liu et al., 2023a; b), LLaVA-1.6 (Liu et al., 2024), Tinny-LLaVA (Zhou et al., 2024), MiniCPM (Hu et al., 2024), Idefics-2  [^1] https://huggingface.co/blog/idefics2, Deepseek-VL(Lu et al., 2024), CogVLM (Wang et al., 2023; Hong et al., 2023), XComposer-4KHD (Zhang et al., 2023), Mantis (Jiang et al., 2024), InternVL-1.5 and InternVL-2 (Chen et al., 2023; 2024) as our MLLMs baseline, and LLaMA3, Yi, Yi-1.5 (AI et al., 2024), Qwen-1.5 (Bai et al., 2023a) and Qwen2 (Yang et al., 2024) as LLMs baseline.
-To ensure fairness in the evaluations, we use the unified evaluation framework provided by MLLMs-eval to conduct evaluation experiments on LIME-M.
+We select LLaVA-1.5 (Liu et al., 2023a; b), LLaVA-1.6 (Liu et al., 2024), Tinny-LLaVA (Zhou et al., 2024), MiniCPM (Hu et al., 2024), Idefics-2  [^1] https://huggingface.co/blog/idefics2, Deepseek-VL(Lu et al., 2024), CogVLM (Wang et al., 2023; Hong et al., 2023), XComposer-4KHD (Zhang et al., 2023), Mantis (Jiang et al., 2024), InternVL-1.5 and InternVL-2 (Chen et al., 2023; 2024b) as our MLLMs baseline, and LLaMA3, Yi, Yi-1.5 (AI et al., 2024), Qwen-1.5 (Bai et al., 2023a) and Qwen2 (Yang et al., 2024) as LLMs baseline.
+To ensure fairness in the evaluations, we use the unified evaluation framework provided by MLLMs-eval to conduct evaluation experiments on LIME.
 For models not supported by this framework, we refine the inference code provided by the model developers to make it compatible with the new models for the sake of aligning the results of different models.
 Metrics
-For most tasks included in LIME-M, we reference the metrics computation methods used in MLLMs-eval.
+For most tasks included in LIME, we reference the metrics computation methods used in MLLMs-eval.
 Specifically, for tasks such as AI2D, ScienceQA, OCRBench, and POPE, we calculate the accuracy of the extracted responses.
@@
 5.1 Main Result
-T/F Common VQA InfoVQA ScienceQA OCR Captioning
-Model Size Overall POPE $\uparrow$ TVQA $\uparrow$ OK VQA $\uparrow$ ChQA $\uparrow$ IVQA $\uparrow$ AI2D $\uparrow$ SciQA $\uparrow$ OCRBen $\uparrow$ COCO $\uparrow$ TCaps $\uparrow$
-InternVL-2 2023 ( - ) 40B 64.19 51.69 77.98 19.45 88.33 73.92 69.20 98.33 34.90 63.10 48.94
-Qwen2-VL 2023b ($\uparrow$ 1) 7B 62.56 53.05 74.56 28.18 83.17 78.50 58.20 75.00 35.70 68.74 110.60
-InternVL-1.5 2024 ($\downarrow$ 1) 26B 62.00 51.69 69.88 35.47 87.00 70.92 54.81 91.67 27.90 69.24 112.63
-InternVL-2 2023 ( - ) 26B 61.48 54.63 75.20 16.08 87.67 70.58 62.80 96.33 32.70 76.18 56.91
-LLaVA-OneVision 2024 ( - ) 7B 60.28 52.37 65.22 37.32 80.83 68.17 59.20 92.00 22.00 104.74 108.18
-InternVL-2 2023 ( - ) 8B 59.52 49.21 66.10 24.20 83.75 68.25 60.40 95.67 32.50 42.55 25.44
-XComposer2-4KHD 2023 ($\uparrow$ 4) 7B 55.67 46.28 60.30 28.13 80.42 66.17 54.90 70.00 24.40 97.07 78.07
-InternVL-2 2023($\downarrow$ 1) 4B 54.86 47.18 62.29 17.48 81.92 60.50 54.00 94.33 30.90 35.99 21.67
-CogVLM-2 2024 ($\uparrow$ 6) 19B 52.02 51.02 69.46 4.92 80.33 59.50 45.00 84.00 31.40 23.67 34.01
-Qwen2-VL 2023b ($\uparrow$ 5) 2B 51.59 50.79 70.70 16.87 67.50 65.00 42.90 58.00 31.30 75.06 101.72
-InternVL-2 2023 ($\downarrow$ 2) 2B 51.25 50.79 59.56 21.87 71.75 54.00 45.80 92.33 31.00 51.95 42.59
-CogVLM-1 2023 ($\uparrow$ 1) 17B 49.53 55.10 71.20 31.70 61.67 57.25 31.40 53.67 19.30 29.28 38.56
-Cambrian 2024 ($\downarrow$ 5) 34B 48.78 49.44 57.28 22.03 71.83 43.17 54.80 78.33 18.00 4.27 4.97
-Cambrian 2024 ($\downarrow$ 4) 13B 47.06 50.79 58.93 24.13 69.25 42.83 45.20 62.67 19.50 7.40 6.52
-Cambrian 2024 ($\downarrow$ 3) 8B 46.41 49.89 59.00 25.23 69.42 37.67 43.60 69.00 19.80 5.85 6.41
-InternVL-2 2023 ($\uparrow$ 2) 1B 45.88 52.82 57.55 15.37 65.83 46.25 37.00 84.33 29.90 15.19 13.19
-LLaVA-1.6 2024 ($\uparrow$ 3) 34B 42.73 47.00 51.20 10.40 64.33 42.08 49.60 64.67 17.10 84.25 48.25
-MiniCPM-LLaMA3-2.5 2024 ($\downarrow$ 3) 8B 40.62 43.10 61.80 25.30 69.00 48.10 6.60 6.60 25.70 31.91 39.86
-LLaVA-OneVision 2024 ($\uparrow$ 4) 0.5B 39.85 48.98 48.61 23.13 55.42 40.67 31.20 53.00 19.70 96.40 90.28
-LLaVA-LLaMA3 2023 ($\downarrow$ 3) 8B 39.81 44.24 40.01 34.72 64.42 22.25 40.72 61.67 13.90 99.35 48.71
-Mantis-Idefics-2 2024 ( - ) 8B 38.09 44.24 44.51 29.07 59.00 20.50 35.80 70.00 14.80 61.82 103.07
-Deepseek-VL 2024 ($\uparrow$ 2) 7B 37.17 48.50 44.80 25.00 54.67 22.33 37.00 68.33 11.70 54.22 83.21
-LLaVA-1.6-vicuna 2024 ($\downarrow$ 4) 13B 35.93 43.10 43.90 16.10 54.50 28.75 38.10 53.00 14.70 76.62 47.83
-Idefics-2 2024 ($\downarrow$ 2) 8B 34.85 42.00 56.50 35.60 13.08 23.92 38.10 78.67 19.60 61.23 94.51
-LLaVA-1.6-vicuna 2024 ( - ) 7B 29.27 41.10 39.00 12.50 43.08 22.67 27.10 47.33 10.90 76.05 48.35
-Mantis-SigLIP 2024 ($\uparrow$ 1) 8B 28.70 45.60 26.34 32.43 35.33 16.25 27.70 62.67 4.90 68.16 81.21
-MiniCPM 2024 ($\uparrow$ 2) 1.0 25.58 44.00 37.00 6.20 35.75 13.42 27.90 60.67 6.70 68.96 76.65
-LLaVA-1.5 2023a ($\downarrow$ 2) 13B 20.12 36.60 19.50 32.10 5.50 12.42 25.90 48.33 2.70 80.89 68.73
-LLaVA-1.5 2023a ($\downarrow$ 1) 7B 16.98 32.51 16.50 23.43 5.25 9.08 24.05 49.00 2.20 79.20 65.73
-InstructBLIP-vicuna 2023 ( - ) 7B 15.30 45.10 11.40 22.10 3.00 9.08 21.90 34.33 2.00 102.08 53.14
-Tiny-LLaVA-1 2024 ( - ) 1.4B 13.76 37.00 18.70 0.90 4.50 12.17 22.80 44.67 1.60 63.19 58.91
-InternVL-2 40B 80.31 89.23 82.59 50.98 85.52 76.08 85.88 98.56 79.90 99.15 62.03
-InternVL-1.5 26B 79.49 88.90 79.00 60.70 83.70 72.50 78.90 94.50 71.40 95.80 148.10
-Qwen2-VL 7B 79.14 88.17 80.92 55.68 83.32 78.86 80.73 85.57 81.20 92.13 144.36
-InternVL-2 26B 78.82 88.64 82.06 48.50 84.44 72.72 83.16 97.47 77.60 110.30 80.10
-LLaVA-OneVision 7B 78.71 89.17 76.02 60.98 80.12 70.69 81.38 95.88 62.10 140.45 136.97
-InternVL-2 8B 77.84 87.90 77.00 52.02 82.48 70.65 82.25 97.03 76.50 89.77 36.70
-InternVL-2 4B 73.97 87.71 74.51 38.43 81.04 65.19 78.08 96.03 75.00 54.08 30.17
-Cambrian 34B 73.26 88.46 72.11 52.07 74.60 51.48 80.41 85.52 59.00 8.18 6.08
-InternVL-2 2B 73.00 88.90 72.39 43.74 74.72 57.69 72.70 94.25 75.50 79.52 59.81
-Cambrian 13B 72.39 88.53 73.07 53.28 72.60 50.73 73.93 79.08 61.40 14.33 9.44
-XComposer-4KHD 7B 71.93 87.00 74.30 51.90 80.60 72.80 34.40 96.00 66.90 134.00 111.40
-Cambrian 8B 71.84 88.24 72.47 52.17 73.44 48.05 72.99 80.32 61.60 9.13 7.97
-CogVLM-1 17B 71.34 88.90 79.70 46.90 67.00 63.30 61.90 70.50 59.10 28.40 44.70
-MiniCPM-LLaMA3-2.5 8B 71.22 88.00 75.00 52.30 72.90 56.90 71.70 53.00 69.40 35.50 52.90
-Qwen2-VL 2B 70.86 87.78 78.70 40.59 73.12 67.15 70.21 77.89 75.30 103.52 131.92
-CogVLM-2 19B 69.93 87.56 77.59 18.51 79.84 62.62 72.41 90.93 76.60 24.10 42.23
-LLaVA-LLaMA3 8B 69.74 87.80 65.40 60.20 69.30 37.60 71.60 73.30 55.00 135.00 69.60
-InternVL-2 1B 68.46 87.94 69.67 33.84 71.40 52.02 62.56 89.59 74.20 49.34 18.03
-LLaVA-1.6-vicuna 13B 67.29 87.50 67.00 46.30 62.20 41.50 70.40 73.50 55.00 101.90 67.30
-LLaVA-1.6 34B 67.22 85.60 68.90 31.00 67.40 51.90 76.10 82.70 58.60 114.40 69.10
-Mantis-Idefics-2 8B 66.91 86.90 63.51 52.50 63.56 31.17 66.81 81.80 54.20 79.42 134.08
-Idefics-2 8B 66.73 86.80 71.30 53.90 26.40 37.00 69.20 87.20 61.60 71.90 119.10
-LLaVA-OneVision 0.5B 65.65 88.33 65.85 44.17 61.36 46.23 57.09 67.03 57.60 131.90 120.81
-Deepseek-VL 7B 65.62 87.10 63.20 48.70 60.60 34.30 63.40 81.70 43.30 67.60 110.10
-LLaVA-1.6-vicuna 7B 64.80 87.60 64.90 44.20 55.00 37.00 65.30 70.20 52.40 100.00 72.00
-LLaVA-1.5 13B 59.58 87.10 48.70 58.30 18.10 29.50 59.40 72.80 33.60 115.40 104.00
-Mantis-SigLIP 8B 58.96 81.47 49.59 52.90 42.56 26.56 57.84 75.36 34.50 91.37 111.43
-LLaVA-1.5 7B 57.27 87.00 46.10 53.40 18.20 25.80 55.20 69.50 31.50 109.00 98.00
-MiniCPM 1.0 56.18 85.10 55.30 47.30 15.40 20.10 56.90 43.00 60.00 25.90 41.60
-InstructBLIP-vicuna 7B 47.87 85.00 33.20 45.20 12.50 22.90 34.00 36.40 25.90 141.40 74.00
-Tiny-LLaVA-1 1.4B 34.30 56.30 38.50 3.80 11.10 22.20 32.30 58.20 17.20 80.90 83.10
-Table 2: Comparing overall scores of LIME-M and Original.
-Top: results on LIME-M, Bottom: results on the original dataset.
-The arrow next to the model name indicates the change in ranking on LIME-M compared to the original dataset.
+Model Size LIME Original Reasoning VQA InfoQA SciQA OCR Caption
+InternVL-2 2023 40B 64.15 ( - ) 80.31 51.69 48.72 81.12 75.92 34.90 56.02
+Qwen2-VL 2023b 7B 62.53 ($\uparrow$ 1) 79.14 53.05 51.37 80.83 62.08 35.70 89.67
+InternVL-1.5 2024b 26B 61.97 ($\downarrow$ 1) 79.49 51.69 52.68 78.96 63.32 27.90 90.93
+InternVL-2 2023 26B 61.46 ( - ) 78.82 54.63 45.64 79.12 70.54 32.70 66.54
+LLaVA-OneVision 2024 7B 60.25 ( - ) 78.71 52.37 51.27 74.50 66.77 22.00 106.46
+InternVL-2 2023 8B 59.49 ( - ) 77.84 49.21 45.15 76.00 68.54 32.50 34.00
+XComposer2-4KHD 2023 7B 55.64 ($\uparrow$ 4) 71.93 46.28 44.22 73.29 58.38 24.40 87.57
+InternVL-2 2023 4B 54.84 ($\downarrow$ 1) 73.97 47.18 39.89 71.21 63.31 30.90 28.83
+CogVLM-2 2024 19B 52.02 ($\uparrow$ 6) 69.93 51.02 37.19 69.92 54.00 31.40 28.84
+Qwen2-VL 2023b 2B 51.59 ($\uparrow$ 5) 70.86 50.79 43.78 66.25 46.38 31.30 88.39
+InternVL-2 2023 2B 51.25 ($\downarrow$ 2) 73.00 50.79 40.71 62.88 56.54 31.00 47.27
+CogVLM-1 2023 17B 49.55 ($\uparrow$ 1) 71.34 55.10 51.45 59.46 36.54 19.30 33.92
+Cambrian 2024 34B 48.79 ($\downarrow$ 5) 73.26 49.44 39.66 57.50 60.23 18.00 4.62
+Cambrian 2024 13B 47.07 ($\downarrow$ 4) 72.39 50.79 41.53 56.04 49.23 19.50 6.96
+Cambrian 2024 8B 46.42 ($\downarrow$ 3) 71.84 49.89 42.12 53.55 49.46 19.80 6.13
+InternVL-2 2023 1B 45.90 ($\uparrow$ 2) 68.46 52.82 36.46 56.04 47.92 29.90 14.19
+LLaVA-1.6 2024 34B 42.74 ($\uparrow$ 3) 67.22 47.00 30.80 53.21 53.08 17.10 66.25
+MiniCPM-LLaMA3-2.5 2024 8B 40.63 ($\downarrow$ 3) 71.22 43.10 43.55 58.55 6.60 25.70 35.89
+LLaVA-OneVision 2024 0.5B 39.88 ($\uparrow$ 4) 65.65 48.98 35.87 48.04 36.23 19.70 93.34
+LLaVA-LLaMA3 2023 8B 39.82 ($\downarrow$ 3) 69.74 44.24 37.36 43.33 45.56 13.90 74.03
+Mantis-Idefics-2 2024 8B 38.11 ( - ) 66.91 44.24 36.79 39.75 43.69 14.80 82.44
+Deepseek-VL 2024 7B 37.20 ($\uparrow$ 2) 65.62 48.50 34.90 38.50 44.23 11.70 68.72
+LLaVA-1.6-vicuna 2024 13B 35.95 ($\downarrow$ 4) 67.29 43.10 30.00 41.63 41.54 14.70 62.23
+Idefics-2 2024 8B 34.88 ($\downarrow$ 2) 66.73 42.00 46.05 18.50 47.46 19.60 77.87
+LLaVA-1.6-vicuna 2024 7B 29.30 ( - ) 64.80 41.10 25.75 32.88 31.77 10.90 62.20
+Mantis-SigLIP 2024 8B 28.76 ($\uparrow$ 1) 58.96 45.60 29.39 25.79 35.77 4.90 74.69
+MiniCPM 2024 1.0 25.63 ($\uparrow$ 2) 56.18 44.00 21.60 24.58 35.46 6.70 72.80
+LLaVA-1.5 2023a 13B 20.17 ($\downarrow$ 2) 59.58 36.60 25.80 8.96 31.08 2.70 74.81
+LLaVA-1.5 2023a 7B 17.03 ($\downarrow$ 1) 57.27 32.51 19.97 7.17 29.81 2.20 72.47
+InstructBLIP-vicuna 2023 7B 15.39 ( - ) 47.87 45.10 16.75 6.04 24.77 2.00 77.61
+Tiny-LLaVA-1 2024 1.4B 13.83 ( - ) 34.30 37.00 9.80 8.33 27.85 1.60 61.05
+Table 2: Comparing overall scores of LIME and Original.
+The arrow next to the LIME score indicates the change in ranking on LIME compared to the original dataset.
 $\uparrow$: upward shift, $\downarrow$: downward shift, and -: no change.
-As shown in Tab 2, we evaluate the mainstream MLLMs on our LIME-M benchmark.
+As shown in Tab 2, we evaluate the mainstream MLLMs on our LIME benchmark.
 Overall, models with larger parameter sizes and newer versions tend to have higher overall scores.
@@
 The performance of InternVL-2-Small(1B-8B), CogVLM series, and Cambrian series are second only to InternVL-2-Large and LLaVA-OneVision-7B, with their overall scores ranging from 45% to 60%.
-Comparing the overall score of LIME-M and Origin benchmarks, it can be observed that certain model series, such as Cambrian and LLaVA-1.5, their overall scores experience a decline.
+Comparing the overall score of LIME and Origin benchmarks, it can be observed that certain model series, such as Cambrian and LLaVA-1.5, their overall scores experience a decline.
 Conversely, the CogVLM and LLaVA-OneVision series’ overall scores show an improvement.
@@
 We believe this limitation may be due to the constraints of the language model component of MLLMs.
-Figure 4: Correlation distribution between LIME-M and Wildvison Elo.
+Figure 4: Correlation distribution between LIME and Wildvison Elo.
 5.2 Correlation Analysis
-Figure 4 illustrates the correlation between the various sub-tasks in LIME-M and WildVision Bench.
-Most tasks in LIME-M exhibit a strong positive correlation with WildVision Bench.
+Figure 4 illustrates the correlation between the various sub-tasks in LIME and WildVision Bench.
+Most tasks in LIME exhibit a strong positive correlation with WildVision Bench.
 There are six subtasks that have correlation scores exceeding 80%.
-Additionally, the overall score of LIME-M correlates at 91% with WV-Elo, which is higher than any individual sub-task and the original bench’s correlates, demonstrating that the overall score of LIME-M provides a more comprehensive reflection of MLLMs capabilities.
+Additionally, the overall score of LIME correlates at 91% with WV-Elo, which is higher than any individual sub-task and the original bench’s correlates, demonstrating that the overall score of LIME provides a more comprehensive reflection of MLLMs capabilities.
 Automated evaluation metrics (e.g., CIDEr) cannot effectively assess the performance of LVMs in captioning tasks.
@@
 Consequently, we exclude it from the overall score calculation.
-There is a certain degree of correlation between the sub-tasks in LIME-M.
+There is a certain degree of correlation between the sub-tasks in LIME.
 On the one hand, the relevance of TextVQA, InfographicVQA and OCRBench are relatively high.
@@
 We assume that ChartQA and infoVQA may also necessitate the use of additional common knowledge by the models to solve problems.
-5.3 Effectiveness of LIME-M
+5.3 Effectiveness of LIME
 Figure 5: with the same series of models, the distribution differences of various Parameter sizes. Left($\bigstar$): LLaVA-1.6 series, Right($\blacktriangle$): InternVL-2 series
 Model series Dataset GiNi stdev
-InternVL-2 LIME-M Original 0.063 0.030 6.864 4.421
-Cambrian LIME-M Original 0.006 0.002 1.227 0.715
-LLaVA-1.6 LIME-M Original 0.042 0.004 6.730 1.418
+InternVL-2 LIME Original 0.063 0.030 6.864 4.421
+Cambrian LIME Original 0.006 0.002 1.227 0.715
+LLaVA-1.6 LIME Original 0.042 0.004 6.730 1.418
 Table 3: Statistics on the score distributions across different model series
 Model size Dataset GiNi stdev
-7B LIME-M Original 0.271 0.086 19.041 10.836
-8B LIME-M Original 0.128 0.046 10.685 6.270
-13B LIME-M Original 0.174 0.043 13.536 6.446
+7B LIME Original 0.271 0.086 19.041 10.836
+8B LIME Original 0.128 0.046 10.685 6.270
+13B LIME Original 0.174 0.043 13.536 6.446
 Table 4: Statistics on the score distributions across different model sizes.
-LIME-M provides a more challenging evaluation for MLLMs.
-As shown in Tab 2, The MLLMs’ performances on the LIME-M are less than that on Original Bench on most tasks.
-Compared to the Origin benchmark, different MLLMs show a larger score range on our LIME-M, indicating that our LIME-M can better reflect the performance differences between models with a smaller amount of data.
+LIME provides a more challenging evaluation for MLLMs.
+As shown in Tab 2, The MLLMs’ performances on the LIME are less than that on Original Bench on most tasks.
+Compared to the Origin benchmark, different MLLMs show a larger score range on our LIME, indicating that our LIME can better reflect the performance differences between models with a smaller amount of data.
 Furthermore, we compare the score variations across different model series and model sizes.
 Figure 5 illustrates a clear positive correlation between model performance and model size within the same model series.
-Notably, LIME-M exhibits a more dispersed score distribution, effectively highlighting the differences in model performance.
+Notably, LIME exhibits a more dispersed score distribution, effectively highlighting the differences in model performance.
 In Table 5.3 and 5.3, the Gini coefficient and standard deviation are used to measure the differences in overall score distribution across same model series and model sizes.
 The larger the Gini coefficient and standard deviation, the greater the disparity in data distribution.
-It can be observed that, whether within the same model series or the same model size, LIME-M achieves higher Gini and standard deviation values compared to the original bench.
-This indicates that LIME-M can better differentiate the performance differences between various models.
+It can be observed that, whether within the same model series or the same model size, LIME achieves higher Gini and standard deviation values compared to the original bench.
+This indicates that LIME can better differentiate the performance differences between various models.
 AI2D ScienceQA
-Model LIME-M Original LIME-M Original
+Model LIME Original LIME Original
 LLaMA3-8B 18.10 46.76 33.33 59.35
@@
 Yi-1.5-34B-Chat 25.20 60.69 46.00 70.55
-Figure 6: Comparing text only results of LIME-M and original bench. Left: text only results between LIME-M and Original on AI2D and ScienceQA; Right: average score comparison of Original and LIME-M.
-LIME-M eliminates potential data leakage.
+Figure 6: Comparing text only results of LIME and original bench. Left: text only results between LIME and Original on AI2D and ScienceQA; Right: average score comparison of Original and LIME.
+LIME eliminates potential data leakage.
 For multimodal question answering tasks, visual information input is essential, and LLMs are unable to provide correct answers due to they cannot perceive the content within the image.
@@
 The average score for AI2D is close to 55%, and for ScienceQA, it exceeds 60%, which shows that data from AI2D and ScienceQA in Original are highly likely to have been exposed to the training data of LLMs.
-In contrast, the LIME-M has eliminated this potential threat, achieving scores below 25% in AI2D and close to 40% in ScienceQA.
+In contrast, the LIME has eliminated this potential threat, achieving scores below 25% in AI2D and close to 40% in ScienceQA.
 According to the results in Figure 6 (left), LLMs tend to score higher on ScienceQA than on AI2D.
@@
 Qwen2-72B 26.3 6.9 2.7 10.8 4.4 10.6 36.3 45.2 5.2 16.8
-VB Yi-1.5-9B-Chat 22.1 2.3 0.3 3.1 0.0 7.8 40.0 0.0 0.2 5.8
+LIME Yi-1.5-9B-Chat 22.1 2.3 0.3 3.1 0.0 7.8 40.0 0.0 0.2 5.8
 LLaMA3-8B 49.0 11.4 3.1 18.6 19.3 32.5 46.9 59.5 6.5 26.4
@@
 Qwen2-72B 60.4 10.5 3.5 15.7 20.5 24.2 34.3 67.9 6.8 28.7
-OB Yi-1.5-9B-Chat 24.7 3.1 0.0 5.9 0.5 7.8 32.7 31.7 0.2 5.8
+Original Yi-1.5-9B-Chat 24.7 3.1 0.0 5.9 0.5 7.8 32.7 31.7 0.2 5.8
 Table 5: Text-only with VD results: With the condition of providing only text QA information and VD information, the performance comparison between vlms-bench and origin bench.
@@
 As shown in Tab 5, to investigate whether the remaining samples need to be answered by using textual and image information, we conducted experiments using LLMs to generate answers on both the Original Benchmark and MLLMs Benchmark under QVD (question + visual description) setting.
-LIME-M requires MLLMs to perceive deeper levels of image information.
-Especially in tasks such as AI2D, OCRBench, and TCaps, the scores of LLMs on LIME-M are significantly lower than on the Original Benchmark when provided with only the questions and simple image descriptions.
-This indicates that, after removing some of the simpler questions, LIME-M is better at testing the models’ ability to perceive image details.
+LIME requires MLLMs to perceive deeper levels of image information.
+Especially in tasks such as AI2D, OCRBench, and TCaps, the scores of LLMs on LIME are significantly lower than on the Original Benchmark when provided with only the questions and simple image descriptions.
+This indicates that, after removing some of the simpler questions, LIME is better at testing the models’ ability to perceive image details.
 6 Related work
@@
 Consequently, numerous more difficult and diverse evaluation benchmarks have emerged in recent years to assess the capabilities of MLLMs comprehensively.
-For instance, MMMU (Yue et al., 2024) and CMMMU (Zhang et al., 2024), are comprehensive benchmark tests for university-level multidisciplinary multimodal understanding and reasoning, spanning various disciplines and fields such as Arts and Design, Business, Science, Health Medicine, Humanities and Social Sciences, and Technology Engineering.
-MMBench (Liu et al., 2023c) has developed a comprehensive evaluation pipeline that offers fine-grained capability assessment and robust evaluation metrics.
+For instance, MMMU (Yue et al., 2024) and CMMMU (Zhang et al., 2024), are comprehensive benchmark tests for university-level multidisciplinary multimodal understanding and reasoning, MMBench (Liu et al., 2023c) has developed a comprehensive evaluation pipeline that offers fine-grained capability assessment and robust evaluation metrics.
 MMRA (Wu et al., 2024b) systematically establishes an association relation system among images to access the multi-image relation mining ability of MLLMs.
@@
 Furthermore, training datasets across different models may contain the samples of those benchmarks, which results in data leakage issues (Fu et al., 2023).
-Therefore, we propose LIME-M: a comprehensive MLLMs benchmark.
-We have carefully selected six types of tasks from the existing mainstream benchmarks and scaled down the existing benchmarks according to clear guidelines.
-It is a streamlined version that retains the essential elements of mainstream MLLMs benchmarks.
+Mmstar (Chen et al., 2024a) and MMLU_Redux (Gema et al., 2024) have identified several issues within current benchmarks.
+Mmstar (Chen et al., 2024a) proposes an automated pipeline to filter benchmark data, aiming to detect potential data leakage, while MMLU_Redux (Gema et al., 2024) focuses on correcting annotation errors.
+However, there is still a pressing need for a comprehensive pipeline that fully addresses the challenges posed by multimodal datasets.
+In response to this, we introduce LIME: LESS IS MORE FOR MLLM EVALUATION.
+We have carefully selected six task types from existing mainstream benchmarks and scaled them down according to clear guidelines.
+This streamlined version retains the core elements of mainstream MLLM benchmarks, providing a more efficient and focused evaluation.
 7 Conclusion
 As MLLMs continue to advance, a notable absence of convenient and high-quality multimodal benchmarks has emerged.
-In response to this, we propose a pipeline aimed at semi-automatically refining existing benchmarks to enhance their quality, culminating in the development of LIME-M, which comprises 9,400 evaluation samples across 6 types of tasks and 10 different benchmark datasets.
-By refining the original benchmarks to filter question difficulty and eliminate potentially problematic items, LIME-M offers a more rigorous evaluation for MLLMs, necessitating a deeper understanding of image information.
-The outcomes of our evaluation experiments demonstrate the heightened challenge posed by LIME-M for MLLMs.
-We anticipate that our approach will contribute to the advancement of MLLM evaluation systems, and we are committed to continually enriching LIME-M with an expanded array of datasets through regular updates and expansions.
+In response to this, we propose a pipeline aimed at semi-automatically refining existing benchmarks to enhance their quality, culminating in the development of LIME, which comprises 9,400 evaluation samples across 6 types of tasks and 10 different benchmark datasets.
+By refining the original benchmarks to filter question difficulty and eliminate potentially problematic items, LIME offers a more rigorous evaluation for MLLMs, necessitating a deeper understanding of image information.
+The outcomes of our evaluation experiments demonstrate the heightened challenge posed by LIME for MLLMs.
+We anticipate that our approach will contribute to the advancement of MLLM evaluation systems, and we are committed to continually enriching LIME with an expanded array of datasets through regular updates and expansions.
 Our ultimate goal is to provide the community with a simpler, more efficient, and more accurate evaluation method and suite for MLLMs.
@@
 A.1 overall Data statics
-Figure 7: The overall percentage distribution of LIME-M.
+Figure 7: The overall percentage distribution of LIME.
 Figure 8: Overview of data statics.
@@
 A.3 Alignment With Real World User Queries
-To bridge the gap between LIME-M and real-world users’ query.
-We construct a similarity search system between the LIME-M and the real-world benchmarks(such as wildvision bench, and VibeEval) Different from single-modal similarity search systems, our system leverages multimodal information as query instructions, which contains text and image features information, to explore the effectiveness of the methods and construct the optimal similarity search system, we propose the following different approaches:
+To bridge the gap between LIME and real-world users’ query.
+We construct a similarity search system between the LIME and the real-world benchmarks(such as wildvision bench, and VibeEval) Different from single-modal similarity search systems, our system leverages multimodal information as query instructions, which contains text and image features information, to explore the effectiveness of the methods and construct the optimal similarity search system, we propose the following different approaches:
   • single modal query
@@
 We only keep the top-1 similar data entry based on similarity.
-Figure 10: Correlation coefficient between LIME-M-fit and WildVision bench.
-In Figure 10, we calculate the correlation coefficient between LIME-M-fit obtained through different query methods and real-world benchmarks.
+Figure 10: Correlation coefficient between LIME-fit and WildVision bench.
+In Figure 10, we calculate the correlation coefficient between LIME-fit obtained through different query methods and real-world benchmarks.
 We find that samples retrieved using image queries align better with the WildVision benchmark compared to those retrieved with text queries.